OpenStudy (dumbsearch2):

f(x) = 3x3 + 12x2 + 3x - 18 @jim_thompson5910

OpenStudy (dumbsearch2):

What are the zeros of that function? :) I found 1 as one of them, but I want to be sure if there are more.

jimthompson5910 (jim_thompson5910):

which one did you find?

OpenStudy (diamondboy):

equate the equation to zero and factor

OpenStudy (dumbsearch2):

A. -1 B. 1 C. 2 D. 3

jimthompson5910 (jim_thompson5910):

so what you can do is plug each into the function and see which one(s) give a result of 0

jimthompson5910 (jim_thompson5910):

simple example: g(x) = 2x + 4 x = -2 is a zero or root of g(x) because g(x) = 2x + 4 g(-2) = 2(-2) + 4 ... replace x with -2 g(-2) = 0 we get an output of 0, so the input x = -2 is a root. This corresponds to an x-intercept of (-2,0).

OpenStudy (dumbsearch2):

-2 is not in the answer choices.

jimthompson5910 (jim_thompson5910):

it's my example

OpenStudy (dumbsearch2):

@jim_thompson5910 I tried all the answer choices, none worked for me.

jimthompson5910 (jim_thompson5910):

I plugged x = -2 into my example function. Try the same with your answer choices.

jimthompson5910 (jim_thompson5910):

show me your work so I can see where you're going wrong

jimthompson5910 (jim_thompson5910):

alternatively, you can graph y = 3x^3 + 12x^2 + 3x - 18 in something like desmos https://www.desmos.com/calculator wherever you see the graph going through the x axis is an x-intercept (ie a root). There will be 3 roots for this equation.

OpenStudy (dumbsearch2):

Sorry, I had made a mistake in the calculation- I forgot to consider the exponent.

OpenStudy (dumbsearch2):

So I was correct originally, it is indeed 1? :) @jim_thompson5910 Thanks so much for your help. Is the answer 1?

jimthompson5910 (jim_thompson5910):

yep, the only root in that list of answer choices is x = 1 the other roots are x = -3 and x = -2 (which aren't listed)